How do i measure the thickness of aluminum foil with the mass of 1.47 grams, area of 305 cm and density of 2.6989 grams?
Answer by mananth(16949)

Density = mass/ volume
density = mass/ (area * tk.)
2.6989= 1.47/(305*tk)
(305*tk)= 1.47/2.6989
tk = 1.47/(2.6989*305)
thickness = 0.001785 cm